# On the Robustness of Kernel Goodness-of-Fit Tests

### v4

@@ 4 Robust Kernel Goodness-of-fit Tests for KSD-Ball Uncertainty Sets
4:  Draw $\mathbf{W}^{b}\sim\textrm{Multinomial}(n;1/n,\ldots,1/n)$ and compute bootstrapped sample $D_{\mathbf{W}^{b}}^{2}$ by (3).
@@
We now propose a novel robust KSD test for the setting where the uncertainty set is a KSD-ball. That is, we consider the composite hypotheses
@@
where $\theta\geq 0$ and for $P\in\mathcal{P}(\mathbb{R}^{d})$ with density $p\in\mathcal{C}^{1}$, we define $\mathcal{B}^{\mathrm{KSD}}(P;\theta)\coloneqq\{Q:\;D(Q,P)\leq\theta\}$. The advantage of using a KSD-ball as the uncertainty set is that it lends naturally to a simple, tractable test that is robust to the contamination models we reviewed in Section˜2.3. We first describe our novel test, which we call robust-KSD test, and show that it is quantitatively robust to KSD balls in the sense of Definition˜2. We then discuss in Section˜4.2 how to choose the uncertainty radius $\theta$ to incorporate other contamination models, such as Huber’s contamination models and density-band models.
@@ 4.1 A Robust KSD Test
where $D(\mathbb{X}_{n})$ is the square-root of the V-statistic (2). The test rejects $H^{\mathrm{C}}_{0}$ for large values of $\Delta_{\theta}(\mathbb{X}_{n})$. The decision threshold should be chosen to control the Type-I error rate for all possible $Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, i.e., we require $\gamma=\gamma(\mathbb{X}_{n})$ so that $\sup_{Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\Delta_{n}(\mathbb{X}_{n})>\gamma)\leq\alpha$. One approach to construct $\gamma$ is to use deviation inequalities for bounded functions as done in Gretton et al., (2012, Corollary 11). However, such approach tends to be overly conservative (Gretton et al.,, 2012). Instead, we propose a bootstrap procedure to construct a decision threshold. We will show that choosing $\gamma=q_{\infty,1-\alpha}(\mathbb{X}_{n})$, the square-root of the (population) bootstrap quantile defined in (4), gives the desired Type-I error control asymptotically. Our robust-KSD test therefore rejects $H^{\mathrm{C}}_{0}:Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ if $\Delta_{\theta}(\mathbb{X}_{n})>q_{\infty,1-\alpha}(\mathbb{X}_{n})$. This is summarized in Algorithm˜1.
@@
Remark 4.1 (Pointwise and uniform controls).
The Type-I error control shown in Theorem 4 is pointwise over the null distributions $Q$. For robust tests, a more desirable control is often a uniform one, whereby the supremum over $Q$ is taken before the limit over $n$ (Lehmann and Romano,, 2022, Chapter 11). For the proposed test, a uniform control can be shown if the bootstrapped quantile is replaced by the ground-truth quantile, say $q_{1-\alpha}^{\ast}(\mathbb{X}_{n})$. That is (see also Remark˜20.1 in the appendix),
@@
However, extending this to the bootstrapped quantile requires uniformly bounding the bootstrap approximation errors, which is non-trivial. We leave this as an open question for future research. In the literature, most kernel-based robust tests that offer uniform controls are two-sample tests using either permutation (Schrab and Kim,, 2024) or deviation bounds (Sun and Zou,, 2023). Yet, these approaches are either not applicable to our one-sample setting or too conservative. This is why our test uses bootstrapping. Another kernel robust test that uses bootstrapping is Key et al., (2025), but their Type-I error control is also pointwise. In Appendix˜D, we propose an alternative robust test that leverages deviation bounds in a manner similar to Sun and Zou, (2023) to achieve uniform controls, but at the expense of a lower test power.
@@ 4.2 Choosing the Uncertainty Radius
This result suggests that, when at most a proportion $\epsilon_{0}$ of data is corrupted, setting $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ will ensure quantitative robustness to $\mathcal{P}(P;\epsilon_{0})$. Note that using a similar argument as in the proof of Proposition˜5, Huber’s contamination models can also be related to the MMD balls; see Section˜B.2.
@@
Remark 5.1.
Proposition˜5 holds for Huber’s contamination models of the form $Q=(1-\epsilon)P+\epsilon R$ with $\epsilon\leq\epsilon_{0}$. In particular, it holds for any contamination distribution $R$. This agnosticism can be useful in practice, because the exact form of contamination is often unknown. On the other hand, when prior knowledge about $R$ is available, it is also possible to incorporate it into the proposed test. For example, if $R$ is known to have a support bounded by some $B>0$, then the bound in Proposition˜5 can be tightened by replacing the worst-case bound $\tau_{\infty}=\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})$ with the localized version $\tau_{\infty}=\sup_{\|\mathbf{x}\|_{2}\leq B}u_{p}(\mathbf{x},\mathbf{x})$. As discussed in Fauß et al., (2021), such assumptions are realistic in, e.g., highly regulated experimental environments, where any outliers are known to lie within bounded regions.
@@
Proposition˜5 and 6 also reveal a limitation of our robust KSD test—it can be overly conservative to alternatives that are not the intended contamination but still lie within the chosen KSD-ball uncertainty set. For example, to ensure robustness to Huber’s contamination models (6) with tolerance $\epsilon_{0}$, Proposition˜5 suggests setting the uncertainty radius to $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$. However, the resulting KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ contains not only Huber’s models, but also other distributions such as density-band models. As a result, the robust KSD test will control Type-I errors for all these distributions, regardless of whether they are the intended contamination type. This limitation is not unique to our method; it is a generic drawback of all robust tests based on uncertainty sets (Fauß et al.,, 2021).
@@ 5.1.1 Decay Rate of Weighting Function
We first draw random samples $\mathbb{X}_{n}$ of size $n=500$ from $Q=(1-\epsilon)P+\epsilon\delta_{z}$, for different values of $\epsilon\in[0,1]$ and $z\in\mathbb{R}$. This setting mimics the presence of outliers at $z$. Given a stationary kernel $h$, a natural question is “how does the choice of weight $w$ affect the Stein kernel and the test power?”. On the left-hand side of Figure˜2, we plot the Stein kernel $x\mapsto u_{p}(x,x)$ for different values of $b$. As $b$ grows, the tails of the Stein kernel are progressively down-weighted. We then plot the rejection probability of the standard KSD tests using these Stein kernels when the outlier is $z=10$ on the right-hand side of Figure˜2. As $b$ increases, the rejection probability decreases, suggesting that the test power is lower when the tails of the Stein kernel are overly down-weighted. This is not surprising since the Stein kernel decays faster for larger values of $b$, making the test insensitive to model deviations at the tails. Ideally, $w$ should decay just enough so that $u_{p}$ remains bounded, but not too fast as it would lose power. This highlights the trade-off between robustness and power in the choice of $w$. A similar trade-off has been observed in robust estimation, where an overly decaying $w$ can impede the estimation accuracy (Barp et al.,, 2019, Appendix F.2). Henceforth, we choose $b=1/2$ to balance this trade-off, as it is the smallest value that ensures $u_{p}$ is bounded for sub-Gaussian models.
@@
Figure 2: Rejection probability of robust-KSD with different bandwidths $\lambda$. “med” is the median heuristic. “KSDAgg” is the test of Schrab et al., (2022). The dashed line is $\alpha=0.05$. The vertical line is the maximal proportion of contamination $\epsilon_{0}=0.05$ controlled by robust-KSD.
@@ 5.1.2 Kernel Bandwidth
We now use the same data set to investigate how the choice of the kernel bandwidth $\lambda$ affects the performance of the robust-KSD test. It is well-known that the bandwidth plays a crucial role in kernel-based tests and can considerably affect the ability of the test to detect model disparities (Gretton et al.,, 2012; Ramdas et al.,, 2015; Reddi et al.,, 2015; Schrab et al.,, 2022, 2023). In general, a smaller bandwidth is better at detecting local differences, while a larger bandwidth is more suited for global deviations (Schrab et al.,, 2023). A common strategy for bandwidth selection in kernel-based testing is the median heuristic. Schrab et al., (2022, 2023) also proposed testing frameworks that aggregate the test result with multiple bandwidths to give the final decision, thereby avoiding bandwidth selection and achieving higher test power. A natural question is therefore: “how does the bandwidth affects the robustness of the proposed test?”.
We run the robust-KSD test with various choices of bandwidths $\lambda^{2}\in\Lambda\cup\{\lambda_{\mathrm{med}}^{2}\}$, where $\Lambda=\{0.01,0.1,1,10,100\}$ and the uncertainty radius $\theta$ is set to control at most $\epsilon=0.05$ contaminations in the sample. We also compare it with the KSDAgg test of Schrab et al., (2022) using the same tilted kernel and aggregating over the collection of bandwidths $\Lambda$. Other configurations of KSDAgg follow the recommendations in Schrab et al., (2022, Section 4.2). This experiment is repeated 200 times to reduce randomness, and the results are reported in Figure˜2. For any choice of $\lambda$, the robust-KSD test is able to control the Type-I error when $\epsilon\leq\epsilon_{0}$, where $\epsilon$ is the proportion of contamination in the data, suggesting that robust-KSD remains well-calibrated for all bandwidths. If instead $\epsilon>\epsilon_{0}$, all robust-KSD tests eventually reject $H^{\mathrm{C}}_{0}$, with $\lambda$ chosen by median heuristic achieving the highest power. The KSDAgg test is not robust and more sensitive to contamination than all robust-KSD tests at all contamination levels $\epsilon$. This is not surprising given that KSDAgg is designed to maximize the test power over multiple bandwidths, rendering this test more sensitive to a wide range of alternatives, including contaminated models. As a result, we henceforth do not include KSDAgg in our experiments. A similar study for the standard KSD test is given in Section˜B.4.
@@ 5.1.3 Standard versus Robust KSD Tests
We also compared our tests with two existing robust kernel tests, namely the robust MMD tests of Schrab and Kim, (2024) (denoted dcMMD) and of Sun and Zou, (2023, Eq. (38)) (denoted MMD-Dev). Both tests are provably quantitatively robust to Huber’s contamination models, but they are two-sample tests that require extra samples from the model $P$. When the same number of samples from $Q$ and from $P$ are used and the cost of simulation is negligible, the cost of a single evaluation of MMD comparable to that of KSD. However, this cost could be prohibitively expensive when $P$ is a more complex model. Implementation details of these tests are deferred to Section˜B.2. As shown in Figure˜3, both tests are more conservative than our proposed KSD tests. The conservativeness of dcMMD is because it is designed for a contamination model that is stronger than Huber’s model, while the conservativeness of MMD-Dev is because it uses a deviation inequality to construct the decision threshold, which is known to be conservative; see, e.g., Gretton et al., (2012).
@@ 5.2 Gaussian-Bernoulli Restricted Boltzmann Machines
Our next experiment is an example of energy-based model called a Gaussian-Bernoulli Restricted Boltzmann Machine (RBM) model (Cho et al.,, 2013). RBMs have been used in a wide range of scientific applications (see Fischer and Igel,, 2014; Zhang et al.,, 2019, for a review), and they are a common benchmarks for assessing kernel GOF tests (Liu et al.,, 2016; Jitkrittum et al.,, 2017; Schrab et al.,, 2022). RBMs are latent-variable models with joint density $p(\mathbf{x},\mathbf{h})\propto\exp(\frac{1}{2}\mathbf{x}^{\top}B\mathbf{h}+\mathbf{b}^{\top}\mathbf{x}+\mathbf{c}^{\top}\mathbf{h}-\frac{1}{2}\|\mathbf{x}\|_{2}^{2})$, where $\mathbf{x}\in\mathbb{R}^{d}$ is an observable variable, $\mathbf{h}\in\{\pm 1\}^{d^{\prime}}$ is a binary hidden variable with latent dimension $d^{\prime}$, and $B\in\mathbb{R}^{d\times d^{\prime}},\mathbf{b}\in\mathbb{R}^{d}$ and $\mathbf{c}\in\mathbb{R}^{d^{\prime}}$ are model parameters. Computing the normalizing constant of the marginal $p(\mathbf{x})$ requires summing over $2^{d^{\prime}}$ terms, so it becomes intractable when $d^{\prime}$ is large. However, its score function has a closed-form expression $\mathbf{s}_{p}(\mathbf{x})=b-\mathbf{x}+B\mathrm{tanh}(B^{\top}\mathbf{x}+\mathbf{c})$, where $\mathrm{tanh}$ is applied entry-wise. We set $d=50$ and $d^{\prime}=10$, and randomly initialize $B,\mathbf{b},\mathbf{c}$ by sampling each entry independently from a $\mathcal{N}(0,1)$. We then generate $n=500$ samples from the model by block Gibbs sampling following Cho et al., (2013); Jitkrittum et al., (2017), and randomly replace $\epsilon\in[0,1]$ proportions of the data with outliers drawn from $R=\mathcal{N}(0,0.1^{2}I_{d})$; see the left and middle of Figure˜5. As shown by the plots, the outliers are clearly abnormal since they have low density values.
@@
When generating the synthetic data from $P$, we run a Gibbs sampler for $7000$ steps and discard the first $2000$ as burn-in and keep every $10$-th datum to reduce correlation. This takes $58.40$ seconds on average, compared with $0.84$ seconds required to perform the robust-KSD test. As a result, it will clearly not be reasonable to use the robust MMD tests from Sun and Zou, (2023); Schrab et al., (2023) here as simulating a large enough number of samples from $P$ would cost orders of magnitude more than the cost of the GOF test. Moreover, the samples generated through Gibbs sampling are no longer i.i.d., thus violating the assumptions in our theoretical results. Extension to the non-independent case could potentially be made following the approach in Chérief-Abdellatif and Alquier, (2022), which studied the robustness of estimators (instead of GOF tests) under correlated data.

### v5

@@ 4 Robust Kernel Goodness-of-fit Tests for KSD-Ball Uncertainty Sets
4:  Draw $\mathbf{W}^{b}\sim\textrm{Multinomial}(n;1/n,\ldots,1/n)$ and compute bootstrapped sample $D_{\mathbf{W}^{b}}^{2}(\mathbb{X}_{n})$ by (3).
@@
We now propose a novel robust KSD test for the setting where the uncertainty set is a KSD-ball. Given $P\in\mathcal{P}(\mathbb{R}^{d})$ with density $p\in\mathcal{C}^{1}$, we consider the composite hypotheses
@@
where $\theta\geq 0$ and $\mathcal{B}^{\mathrm{KSD}}(P;\theta)\coloneqq\{Q:\;D(Q,P)\leq\theta\}$. The advantage of using a KSD-ball as the uncertainty set is that it lends naturally to a simple, tractable test that is robust to the contamination models we reviewed in Section˜2.3. We first describe our novel test, which we call robust-KSD test, and show that it is quantitatively robust to KSD balls in the sense of Definition˜2. We then discuss in Section˜4.2 how to choose the uncertainty radius $\theta$ to incorporate common contamination models, such as Huber’s contamination models and density-band models.
@@ 4.1 A Robust KSD Test
where $D(\mathbb{X}_{n})$ is the square-root of the V-statistic (2). The test rejects $H^{\mathrm{C}}_{0}$ for large values of $\Delta_{\theta}(\mathbb{X}_{n})$. The decision threshold should be chosen to control the Type-I error rate for all possible $Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, i.e., we require $\gamma=\gamma(\mathbb{X}_{n})$ so that $\sup_{Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\Delta_{n}(\mathbb{X}_{n})>\gamma)\leq\alpha$. One approach to construct $\gamma$ is to use deviation inequalities for bounded functions as done in Gretton et al. (2012, Corollary 11). However, such approach tends to be overly conservative (Gretton et al., 2012). Instead, we propose a bootstrap procedure to construct a decision threshold. We will show that choosing $\gamma=q_{\infty,1-\alpha}(\mathbb{X}_{n})$, the square-root of the (population) bootstrap quantile defined in (4), gives the desired Type-I error control asymptotically. Our robust-KSD test therefore rejects $H^{\mathrm{C}}_{0}:Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ if $\Delta_{\theta}(\mathbb{X}_{n})>q_{\infty,1-\alpha}(\mathbb{X}_{n})$. This is summarized in Algorithm˜1.
@@
Remark 7 (Pointwise and uniform controls).
The Type-I error control shown in Theorem 4 is pointwise over the null distributions $Q$. For robust tests, a more desirable control is often a uniform one, whereby the supremum over $Q$ is taken before the limit over $n$ (Lehmann and Romano, 2022, Chapter 11). For the proposed test, a uniform control can be shown if the bootstrapped quantile is replaced by the ground-truth quantile, say $q_{1-\alpha}^{\ast}(\mathbb{X}_{n})$. That is (see also Remark˜9 in the appendix),
@@
However, extending this to the bootstrapped quantile requires uniformly bounding the bootstrap approximation errors, which is non-trivial. We leave this as an open question for future research. In the literature, most kernel-based robust tests that offer uniform controls are two-sample tests using either permutation (Schrab and Kim, 2024) or deviation bounds (Sun and Zou, 2023). Yet, these approaches are either not applicable to our one-sample setting or too conservative. This is why our test uses bootstrapping. Another kernel robust test that uses bootstrapping is Key et al. (2025), but their Type-I error control is also pointwise. In Appendix˜D, we propose an alternative robust test that leverages deviation bounds in a manner similar to Sun and Zou (2023) to achieve uniform controls, but at the expense of a lower test power.
@@ 4.2 Choosing the Uncertainty Radius
This result suggests that, when at most a proportion $\epsilon_{0}$ of data is corrupted, setting $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ will ensure quantitative robustness to $\mathcal{P}(P;\epsilon_{0})$. Note that using a similar proof, Huber’s contamination models can also be related to the MMD balls; see Lemma˜23 in Section˜B.2.
@@
Remark 8.
Proposition˜5 holds for Huber’s contamination models of the form $Q=(1-\epsilon)P+\epsilon R$ with $\epsilon\leq\epsilon_{0}$. In particular, it holds for any contamination distribution $R$. This agnosticism can be useful in practice, because the exact form of contamination is often unknown. On the other hand, when prior knowledge about $R$ is available, it is also possible to incorporate it into the proposed test. For example, if $R$ is known to have a support bounded by some $B>0$, then the bound in Proposition˜5 can be tightened by replacing the worst-case bound $\tau_{\infty}=\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})$ with the localized version $\tau_{\infty}=\sup_{\|\mathbf{x}\|_{2}\leq B}u_{p}(\mathbf{x},\mathbf{x})$. As discussed in Fauß et al. (2021), such assumptions are realistic in, e.g., highly regulated experimental environments, where any outliers are known to lie within bounded regions.
@@
Proposition˜5 and 6 also reveal a limitation of our robust KSD test—it can be overly conservative to alternatives that are not the intended contamination but still lie within the chosen KSD-ball uncertainty set. For example, to ensure robustness to Huber’s contamination models (6) with tolerance $\epsilon_{0}$, Proposition˜5 suggests setting the uncertainty radius to $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$. However, the resulting KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ contains not only Huber’s models, but also other distributions such as density-band models. As a result, the robust KSD test will control Type-I errors for all these distributions, regardless of whether they are the intended contamination type. This limitation is not unique to our method; it is a generic drawback of all robust tests based on uncertainty sets (Fauß et al., 2021).
@@ 5.1.1 Decay Rate of Weighting Function
We first draw random samples $\mathbb{X}_{n}$ of size $n=500$ from $Q=(1-\epsilon)P+\epsilon\delta_{z}$, for different values of $\epsilon\in[0,1]$ and $z\in\mathbb{R}$. This setting mimics the presence of outliers at $z$. Given a stationary kernel $h$, a natural question is “how does the choice of weight $w$ affect the Stein kernel and the test power?”. On the left-hand side of Figure˜2, we plot the Stein kernel $x\mapsto u_{p}(x,x)$ for different values of $b$. As $b$ grows, the tails of the Stein kernel are progressively down-weighted. We then plot the rejection probability of the standard KSD tests using these Stein kernels when the outlier is $z=10$ on the right-hand side of Figure˜2. As $b$ increases, the rejection probability decreases, suggesting that the test power is lower when the tails of the Stein kernel are overly down-weighted. This is not surprising since the Stein kernel decays faster for larger values of $b$, making the test insensitive to model deviations at the tails. Ideally, $w$ should decay just enough so that $u_{p}$ remains bounded, but not too fast as it would lose power. This highlights the trade-off between robustness and power in the choice of $w$. A similar trade-off has been observed in robust estimation, where an overly decaying $w$ can impede the estimation accuracy (Barp et al., 2019, Appendix F.2). Henceforth, we choose $b=1/2$ to balance this trade-off, as it is the smallest value that ensures $u_{p}$ is bounded for sub-Gaussian models.
@@
Figure 2: Rejection probability of robust-KSD with different bandwidths $\lambda$. “med” is the median heuristic. “KSDAgg” is the test of Schrab et al. (2022). The dashed line is $\alpha=0.05$. The vertical line is the maximal proportion of contamination $\epsilon_{0}=0.05$ controlled by robust-KSD.
@@ 5.1.2 Kernel Bandwidth
We now use the same data set to investigate how the choice of the kernel bandwidth $\lambda$ affects the performance of the robust-KSD test. It is well-known that the bandwidth plays a crucial role in kernel-based tests and can considerably affect the ability of the test to detect model disparities (Gretton et al., 2012; Ramdas et al., 2015; Reddi et al., 2015; Schrab et al., 2022, 2023). In general, a smaller bandwidth is better at detecting local differences, while a larger bandwidth is more suited for global deviations (Schrab et al., 2023). A common strategy for bandwidth selection in kernel-based testing is the median heuristic. Schrab et al. (2022, 2023) also proposed testing frameworks that aggregate the test result with multiple bandwidths to give the final decision, thereby avoiding bandwidth selection and achieving higher test power. A natural question is therefore: “how does the bandwidth affects the robustness of the proposed test?”.
We run the robust-KSD test with various choices of bandwidths $\lambda^{2}\in\Lambda\cup\{\lambda_{\mathrm{med}}^{2}\}$, where $\Lambda=\{0.01,0.1,1,10,100\}$ and the uncertainty radius $\theta$ is set to control at most $\epsilon=0.05$ contaminations in the sample. We also compare it with the KSDAgg test of Schrab et al. (2022) using the same tilted kernel and aggregating over the collection of bandwidths $\Lambda$. Other configurations of KSDAgg follow the recommendations in Schrab et al. (2022, Section 4.2). This experiment is repeated 200 times to reduce randomness, and the results are reported in Figure˜2. For any choice of $\lambda$, the robust-KSD test is able to control the Type-I error when $\epsilon\leq\epsilon_{0}$, where $\epsilon$ is the proportion of contamination in the data, suggesting that robust-KSD remains well-calibrated for all bandwidths. If instead $\epsilon>\epsilon_{0}$, all robust-KSD tests eventually reject $H^{\mathrm{C}}_{0}$, with $\lambda$ chosen by median heuristic achieving the highest power. The KSDAgg test is not robust and more sensitive to contamination than all robust-KSD tests at all contamination levels $\epsilon$. This is not surprising given that KSDAgg is designed to maximize the test power over multiple bandwidths, rendering this test more sensitive to a wide range of alternatives, including contaminated models. As a result, we henceforth do not include KSDAgg in our experiments. A similar study for the standard KSD test is given in Section˜B.4.
@@ 5.1.3 Standard versus Robust KSD Tests
We also compared our tests with two existing robust kernel tests, namely the robust MMD tests of Schrab and Kim (2024) (denoted dcMMD) and of Sun and Zou (2023, Eq. 38) (denoted MMD-Dev). Both tests are provably quantitatively robust to Huber’s contamination models, but they are two-sample tests that require extra samples from the model $P$. When the same number of samples from $Q$ and from $P$ are used and the cost of simulation is negligible, the cost of a single evaluation of MMD comparable to that of KSD. However, this cost could be prohibitively expensive when $P$ is a more complex model. Implementation details of these tests are deferred to Section˜B.2. As shown in Figure˜3, both tests are more conservative than our proposed KSD tests. The conservativeness of dcMMD is because it is designed for a contamination model that is stronger than Huber’s model, while the conservativeness of MMD-Dev is because it uses a deviation inequality to construct the decision threshold, which is known to be conservative; see, e.g., Gretton et al. (2012).
@@ 5.2 Gaussian-Bernoulli Restricted Boltzmann Machines
Our next experiment is an example of energy-based model called a Gaussian-Bernoulli Restricted Boltzmann Machine (RBM) model (Cho et al., 2013). RBMs have been used in a wide range of scientific applications (see Fischer and Igel, 2014; Zhang et al., 2019, for a review), and they are a common benchmarks for assessing kernel GOF tests (Liu et al., 2016; Jitkrittum et al., 2017; Schrab et al., 2022). RBMs are latent-variable models with joint density $p(\mathbf{x},\mathbf{h})\propto\exp(\frac{1}{2}\mathbf{x}^{\top}B\mathbf{h}+\mathbf{b}^{\top}\mathbf{x}+\mathbf{c}^{\top}\mathbf{h}-\frac{1}{2}\|\mathbf{x}\|_{2}^{2})$, where $\mathbf{x}\in\mathbb{R}^{d}$ is an observable variable, $\mathbf{h}\in\{\pm 1\}^{d^{\prime}}$ is a binary hidden variable with latent dimension $d^{\prime}$, and $B\in\mathbb{R}^{d\times d^{\prime}},\mathbf{b}\in\mathbb{R}^{d}$ and $\mathbf{c}\in\mathbb{R}^{d^{\prime}}$ are model parameters. Computing the normalizing constant of the marginal $p(\mathbf{x})$ requires summing over $2^{d^{\prime}}$ terms, so it becomes intractable when $d^{\prime}$ is large. However, its score function has a closed-form expression $\mathbf{s}_{p}(\mathbf{x})=b-\mathbf{x}+B\mathrm{tanh}(B^{\top}\mathbf{x}+\mathbf{c})$, where $\mathrm{tanh}$ is applied entry-wise. We set $d=50$ and $d^{\prime}=10$, and randomly initialize $B,\mathbf{b},\mathbf{c}$ by sampling each entry independently from a $\mathcal{N}(0,1)$. We then generate $n=500$ samples from the model by block Gibbs sampling following Cho et al. (2013); Jitkrittum et al. (2017), and randomly replace $\epsilon\in[0,1]$ proportions of the data with outliers drawn from $R=\mathcal{N}(0,0.1^{2}I_{d})$; see the left and middle of Figure˜5. As shown by the plots, the outliers are clearly abnormal since they have low density values.
@@
When generating the synthetic data from $P$, we run a Gibbs sampler for $7000$ steps and discard the first $2000$ as burn-in and keep every $10$-th datum to reduce correlation. This takes $58.40$ seconds on average, compared with $0.84$ seconds required to perform the robust-KSD test. As a result, it will clearly not be reasonable to use the robust MMD tests from Sun and Zou (2023); Schrab et al. (2023) here as simulating a large enough number of samples from $P$ would cost orders of magnitude more than the cost of the GOF test. Moreover, the samples generated through Gibbs sampling are no longer i.i.d., thus violating the assumptions in our theoretical results. Extension to the non-independent case could potentially be made following the approach in Chérief-Abdellatif and Alquier (2022), which studied the robustness of estimators (instead of GOF tests) under correlated data.